OpenStudy (abbycross167):

can someone please help me fix my math problem? please? I will fan and give a medal!! Question~ write the following inequalities in y = mx + b form. Please show all work. –3x –4y > 12 My answer~ Add 3x to both sides -4y>12+3x Divide both sides by -4 (-4) -4y>12+3x(-4) (-4)/-4=1 3/(-4)= 0.75 = 1y>12+0.75x

OpenStudy (abbycross167):

What my teacher said was wrong~ "Watch your signs"

OpenStudy (anonymous):

Am example \[-1>-2\] but multiply by -1 and \[1< 2.\] But then you went to =, so that's ok, it's just a step late. But then when you reintroduce the < > sign, check you pick the right one. Just use numbers to check.

OpenStudy (anonymous):

*An example

OpenStudy (anonymous):

Oh, and it looks like you forgot to include your 12 went you divided. Needed \[\frac{12+3x}{-4}\]

OpenStudy (anonymous):

\[-3x-4y>12\] \[-4y>12+3x\] \[y<-3+-\frac{3}{4}x\]

OpenStudy (abbycross167):

So is −3x−4y>12 −4y>12+3x y<−3+−34x the first three steps?

OpenStudy (anonymous):

Yep, with more explication maybe. You can always check the > < sign is the right way by subbing in numbers like (1,1) at the beginning and end of the question.

OpenStudy (abbycross167):

oh ok... why does y<−3+−34x have the < when the other step have the >?

OpenStudy (anonymous):

it's the \[-1>-2\]but\[(-1)(-1)<(-2)(-1)\]\[1<2\] thing.

OpenStudy (anonymous):

Try it with numbers. convince yourself.

OpenStudy (abbycross167):

@Redcan So would this be the answer? Add 3x to both sides, -4y>12+3x Divide both sides by -4 -3x-4y>12 -4y>12+3x Y<-3+-34x = 1y>12+0.75x

OpenStudy (abbycross167):

@notshh Could you try to help me please?

OpenStudy (notshh):

it was looking good up until this line: Y<-3+-34x

OpenStudy (abbycross167):

yes ma'am/sir, that one I was a little confused on...

OpenStudy (notshh):

-4y>12+3x after dividing -4 from both sides you should get y < (-3/4)x - 3

OpenStudy (abbycross167):

oh okay so, Add 3x to both sides, -4y>12+3x Divide both sides by -4 -3x-4y>12 -4y>12+3x y<(-3/4)x-3 = 1y>12+0.75x This is right now?

OpenStudy (notshh):

how did you go from y<(-3/4)x-3 to 1y>12+0.75x

OpenStudy (abbycross167):

.... Umm, I'm not sure... I thought that Y<-3+-34x was the only thing wrong with it

OpenStudy (notshh):

it's the same equation

OpenStudy (notshh):

everything from Y<-3+-34x onwards was incorrect

OpenStudy (abbycross167):

oh okay